write the equation of the line that passes through (-3,-4) and ( 2-4) in the form y=mx+b

Answer by TimothyLamb(4379)

Calculate slope: m = (y2 - y1)/(x2 - x1)
---
p1 = (-3, -4)
p2 - (2, -4)
---
m = (-4 - -4)/(2 - -3)
m = 0 / 5
m = 0
---
The line is horizontal (slope = 0).
Its y-intercept (b) is the y-coordinate of any of its points:
b = -4
---
The slope-intercept form of the line: y = mx + b
Is this:
y = -4
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php